This is not about math. This all about physics. But please help me :)
here's my questions:
1. An automobile travels with a speed of 88km/hr. How many km does it travel in 0.05 hr?
3. A truck moves along a straight road at an average speed of 10m/s. What is the distance traveled by the truck in 3.2h?
Please help me. Thank you! God bless!

Answer by lwsshak3(11628)

1. An automobile travels with a speed of 88km/hr. How many km does it travel in 0.05 hr?
speed*travel time=distance
88 km/hr*.05 hr=4.4 km
..
3. A truck moves along a straight road at an average speed of 10m/s. What is the distance traveled by the truck in 3.2h?
1 hr=3600sec
3.2 hr=3.2*3600=11520sec
10 m/s*11520s=115200 m or 115.2 km